OpenStudy (anonymous):

sin and cosine of 290 degrees in radians

OpenStudy (paxpolaris):

?

OpenStudy (paxpolaris):

sin and cosine of an angle is just a number .... (there is no radians)

OpenStudy (anonymous):

i need the exact value ...sin and cosin of 930 degrees

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

290 degrees or 930 ?????

OpenStudy (anonymous):

930

OpenStudy (aum):

Subtract as many 360 degrees as possible from 930. 930 - 360 = 570 570 - 360 = 210 210 degrees is in the third quadrant where both sine and cosine are negative. The reference angle for 210 degrees is 30 degrees. \[ \sin(210) = -\sin(30) = -\frac 12 \\ \cos(210) = -\cos(30) = -\frac{\sqrt{3}}{ 2} \]